# An embedding-based distance for temporal graphs

### v1

@@ Abstract
We define a distance between temporal graphs based on graph embeddings built using time-respecting random walks. We study both the case of matched graphs, when there exists a known relation between the nodes, and the unmatched case, when such a relation is unavailable and the graphs may be of different sizes. We illustrate the interest of our distance definition, using both real and synthetic temporal network data, by showing its ability to discriminate between graphs with different structural and temporal properties. Leveraging state-of-the-art machine learning techniques, we propose an efficient implementation of distance computation that is viable for large-scale temporal graphs.
@@ Introduction
Graphs are ubiquitous mathematical entities formed by a set of nodes and one of edges connecting them [1, 2, 3]. They can model a wide range of interacting systems, such as social [4, 5], technological [6], spatial [7] and biological networks [8] and owe their popularity to their ability to encode the complex relational structure of these systems. Real-world systems, moreover, often have temporal properties that cannot be encoded in static graphs and call for modeling based on time-resolved network representations, known as temporal graphs [9]. Examples of these include transportation [10] and ecological networks [11], face-to-face interactions [12, 13, 14], collaboration networks [15, 16, 17], etc.
@@
Given their ubiquitous use in representing such diverse kinds of systems, it is crucial to be able to compare them. In fact, defining and computing a similarity measure between pairs of graphs [18] underpins many important applications and tasks, including machine learning tasks such as graph classification. However, given graphs’ high dimensionality and structural complexity, many different notions of distance can be devised, capturing different properties of interest. Because of this, many definitions of distance for static graphs have been introduced, as reviewed in Refs. [19, 20, 21, 22, 23]. One of the simplest approaches is the edit distance [24] that counts the number of elementary changes (edge or node removal or addition) needed to transform one graph into another. This method only accounts for local information and suffers from the curse of dimensionality [25]. Approaches involving the global graph structure are based on matrix inversion [26, 27] but generally have a high computational cost, and approximations are needed to achieve satisfactory computational performance. The above methods are designed for graph pairs with known correspondence between their nodes, a case we will refer to as matched graphs. The more general case of unmatched graphs involves computing distances for pairs of graphs for which a mapping between nodes is unavailable, and the graphs may have a different size. In this case, one can extract and compare adequately chosen sets of graph features, as done in Refs. [28, 29] or focus on the eigenvalues of appropriate matrix representations [30, 31, 32, 33]. Also, in this case, the computational complexity is a bottleneck, and approximate heuristics are often required [19].
@@
In the case of temporal graphs, the temporal dimension adds to the complexity of graph comparison. Many of the aforementioned distance measures for static graphs have been used to study temporal graphs [34], typically to compare the state of the graph at different points in time [35, 36, 37]. Here, we study the problem of comparing temporal graphs as a whole. In the temporal case, few notions of distance have been proposed: Ref. [38] defines a method to compare temporal graphs based on a set of features that can be tailored to a specific research question; Ref. [39] addresses our same problem of temporal graph comparison but assumes the graphs to be matched and is computationally inefficient; Ref. [40] only considers the unmatched case and uses the comparison of temporal paths statistics to provide a definition of dissimilarity (rather than distance).
@@
Here, we propose a computationally efficient method to compute a distance metric between pairs of temporal graphs, considering both the matched and unmatched cases. We build our comparison on top of graph embeddings, leveraging ideas from both the local and global approaches mentioned above, and encoding differences at the topological as well as at the temporal levels. Commented Python code for implementing our method is available at github.com/lorenzodallamico/G-DynaDist.
@@
Before delving into the details of our main contribution, let us first set some basic definitions.
@@
A temporal graph $\mathcal{G}(\mathcal{V},\mathcal{E},\mathcal{W})$ is a tuple where $\mathcal{V}$ is a set of $n$ nodes, $\mathcal{E}$ a set of temporal edges between nodes, and $\mathcal{W}$ a set of edge weights. All edges $e\in\mathcal{E}$ have the form $e=(i,j,t)$, representing an instantaneous interaction between $i$ and $j$ at time $t$. $w_{e}$ indicates the positive weight of edge $e$. The graph $\mathcal{G}(\mathcal{V},\mathcal{E},\mathcal{W})$ can be represented with a sequence of weighted adjacency matrices, $\{W_{t}\}_{t=1,\dots,T}$, where $T$ is the number of time points. Each of these matrices has size $n\times n$ and entries $(W_{t})_{ij}=w_{(i,j,t)}$ if $(i,j,t)\in\mathcal{E}$ and $(W_{t})_{ij}=0$ otherwise. In the following, we will refer to these matrices as “snapshots” of the temporal graph at given times, and to $T$ as the number of snapshots.
This particular definition is commonly referred to as snapshot graph [41]. In a more general description of temporal networks, each interaction may be represented with a tuple $(i,j,t,\tau)$ where $\tau\in\mathbb{R}^{+}$ is an interaction duration. We adopt the former notation because it is simpler to handle in the following. The two notations are completely interchangeable whenever the time is discrete and a finite temporal resolution $t_{\rm res}$ is set. In fact, all contacts with a duration within $t_{\rm res}$ are considered to be instantaneous, and $(i,j,t,\tau)$ can be represented by a set of interactions of unitary duration: $\{(i,j,t+\delta)\}_{\delta\in\{0,\dots,\tau-1\}}$. For this reason, Definition 1 is sufficiently general, as it can also describe interactions lasting more than one temporal unit. For an arbitrary time resolution, the weight $w_{e}$ associated with a temporal edge $e$ may be used to encode the fraction of $t_{\rm res}$ the two nodes connected by $e$ spent in interaction.
@@
Let $\mathcal{G}_{1}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{W}_{1})$ and $\mathcal{G}_{2}(\mathcal{V}_{2},\mathcal{E}_{2},\mathcal{W}_{2})$ be two temporal graphs with $n_{1}$ and $n_{2}$, respectively. We say that $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are matched if $n_{1}=n_{2}$ and there exists a known bijective function $\pi:\mathcal{V}_{1}\to\mathcal{V}_{2}$ that maps each node of $\mathcal{G}_{1}$ to a node of $\mathcal{G}_{2}$. The two graphs are otherwise said to be unmatched.
When two graphs have the same size, mathematically, there are always many possible mappings $\pi$ between their nodes, but for the matched approach to be valuable, a known mapping associated with an external notion of node identity across the two graphs is necessary (e.g., nodes corresponding to the same persons in two temporal social networks). An implicit consequence of Definition 2 is that all graph pairs with a different number of nodes are considered to be unmatched. We now state the requirements for our distance measure and illustrate our main result.
Figure 1: Distance computation between two temporal graphs. The inputs to the distance function are two temporal graphs $\mathcal{G}_{1}$ (top, orange nodes) and $\mathcal{G}_{2}$ (bottom, green nodes), generally with a different number of snapshots ($T_{1}=5$, $T_{2}=4$, in the example) and a different number of nodes (here $n_{1}=6$, $n_{2}=5$). Each graph is represented as a matrix $P\in\mathbb{R}^{n\times n}$, according to (1), with entry $(ij)$ encoding the probability that a walker goes from $i$ to $j$ following a time-respecting random walk (the walker’s position for each time point is indicated in blue). The matrices $P_{1,2}$ are then embedded using the EDRep algorithm, mapping them to $X_{1}\in\mathbb{R}^{n_{1}\times d}$ and $X_{2}=\mathbb{R}^{n_{2}\times d}$. Finally, the matched – (3) – and unmatched – (4) – distances are calculated. Notice that a necessary condition to compute $d_{\rm m}$ is that $n_{1}=n_{2}$.
@@
In the following, we define two types of distance functions: $d_{\rm m},d_{\rm u}$ between matched and unmatched graphs, respectively. The inputs to these functions are two temporal graphs, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, with a potentially different number of time points (snapshots) $T_{1}$ and $T_{2}$. In both cases, the distances should fulfill the following standard requirements [42]: i) non-negativity; ii) separation axiom; iii) symmetry; iv) triangle inequality. In addition, both distances should also be able by design to detect differences induced by the permutation of time indices. The matched distance should also differentiate between temporal networks differing only by a permutation of node indices, while the unmatched one should be invariant with respect to such transformation.
Our proposed distances are computed in two steps: i) we generate a temporal graph embedding, given by a matrix of size $(n\times d)$ (with $d$ independent of $T$), that encodes relevant topological and temporal properties by leveraging time-respecting random walks on the temporal network; ii) we define the distance in the embedded space (separately treating the matched and the unmatched cases). The distance computation pipeline is illustrated in Fig. 1, and we now describe the steps outlined above in detail.
@@
Graphs are rich data representations, making it challenging to define simple mathematical operations to manipulate them. Graph embeddings provide an interesting avenue for doing so, as they yield a representation in vector space that preserves some of the graphs’ relevant properties [43, 44, 45, 46]. In the case of temporal graphs, the additional challenge of dealing with their temporal dimension has been tackled in many different ways: Ref. [47] represents time as an additional dimension of an adjacency tensor; in evolutionary spectral clustering [48, 49, 50, 51, 52], a different embedding is obtained for each snapshot, and smoothness conditions are implemented on their temporal evolution; more recent works [53, 54, 55] introduced approaches based on time-respecting random walks, which encode relevant temporal properties that are important for dynamical processes occurring on temporal graphs.
@@
We take the latter type of approach as our starting point to generate temporal graph embeddings that are obtained as the solution to an optimization problem. We define a loss function whose argument is a transition matrix $P$ that is the limiting distribution of time-respecting random walks over the temporal graph $\mathcal{G}$ to be embedded. $T$ being the number of snapshots of $\mathcal{G}$, we consider an ensemble of random walks of length drawn uniformly at random between $1$ and $T$, and starting from a randomly chosen node at time $t=1$; at each time step $t$, the walker, situated, e.g., on node $i$, can either move to a random neighbor of $i$ in the snapshot at time $t$ or stay in place on $i$; if the walker is situated on a node that is not active at time $t$, i.e. which has no neighbors in this snapshot, the walker stays in place with probability one. The graph snapshots on the left of Fig. 1 show examples of such time-respecting random walks on the input graphs.
@@
Specifically, given a temporal graph $\mathcal{G}$, with $W_{t}$ the snapshot adjacency matrices of Definition 1 at times $t=1,\cdots,T$, we define for each snapshot $\hat{L}_{t}=(D_{t}+I_{n})^{-1}(W_{t}+I_{n})$, where $D_{t}$ is the degree matrix $D_{t}={\rm diag}(W_{t}\bm{1}_{n})$ of the snapshot, $I_{n}$ is the identity matrix of size $n$, and $\bm{1}_{n}$ is the $n$-sized unity vector. We then define the matrix $P$ as
$P=\frac{1}{T}\sum_{\tau=1}^{T}\hat{L}_{\tau}\hat{L}_{\tau-1}\cdots\hat{L}_{2}\hat{L}_{1}\,\,,$ (1)
The entries of $P$ (which is a probability matrix) depend on the graph structure, the distribution of interaction durations, the correlation between nodes’ interaction patterns, etc. In particular, they are not invariant under permutations of the time indices. We remark that the random walks used here can be of arbitrary length, hence the matrix $P$ is sensitive to all the time scales of the temporal graph considered.
@@
Given the above probability matrix $P$ and a chosen embedding dimensionality $d$, we now want to generate, for each node $i$ of $\mathcal{G}$, a unitary node embedding vector $\bm{x}_{i}\in\mathbb{R}^{d}$. To this aim, letting $X\in\mathbb{R}^{n\times d}$ be the matrix storing in its rows these embedding vectors, we write the loss function as:
@@
This loss function is the sum of the cross entropies between the distributions in the rows of $P$ and the variational distributions in the rows of $Q(X)$, with the addition of a regularization term. The function $\mathcal{L}_{P}$ requires $\mathcal{O}(n^{2})$ operations to be evaluated, since $Z_{i}$ is computed in $\mathcal{O}(n)$ operations for each $i$. We adopt the recently proposed EDRep algorithm [56] that proposes an efficient approximation of $Z_{i}$, making it possible to optimize $\mathcal{L}_{P}$ in $O(n)$ operations. We refer the reader to the Methods section for further details.
The matrix $X$ obtained by optimizing $\mathcal{L}_{P}$ defines the embedding of the temporal graph $\mathcal{G}$, and we now show how to leverage this embedding procedure to define a distance between temporal graphs.
Defining a distance between embeddings
We want to define a distance between temporal graph embeddings that quantifies the similarity of the original temporal graphs according to the requirements stated above. Notice that the output embedding vectors $\bm{x}$ are defined up to an orthogonal transformation applied to the rows of $X$. In fact, suppose $R\in\mathbb{R}^{d\times d}$ is an orthogonal matrix, and $\tilde{\bm{x}}_{i}=R\bm{x}_{i}$, then $\bm{x}_{i}^{T}\bm{x}_{j}=\bm{\tilde{x}}_{i}^{T}\bm{\tilde{x}}_{j}$ and the loss function of (2) is $\mathcal{L}_{P}(\tilde{X})=\mathcal{L}_{P}(X)$. Hence, our distance must be invariant under orthogonal transformations. Moreover, as discussed above, the distance for unmatched graphs must also be invariant with respect to node permutations.
Given these requirements, we now discuss separately our choices to define the matched and unmatched distances. In the Empirical evaluation section, we show the efficacy of these choices with numerical tests.
@@
Let $X_{1}$ and $X_{2}\in\mathbb{R}^{n\times d}$ be the embedding matrices of two matched temporal graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. To satisfy the invariance requirements discussed above, rather than defining our distance directly in terms of the matrices $X_{1}$ and $X_{2}$, we use the auxiliary matrices $M_{X_{1}}=X_{1}X_{1}^{T}$ and $M_{X_{2}}=X_{2}X_{2}^{T}$ of size $n\times n$. These matrices are by construction invariant with respect to orthogonal transformations of the embedding space and encode the global structure of the corresponding temporal graphs (as the embeddings are generated exploiting the whole networks’ structures). We define the distance $d_{\rm m}$ between matched temporal graphs as:
@@
where $\|\cdot\|_{\rm F}$ is the Frobenius norm. Each matrix $M$ encodes the similarity in the embedding space between all pairs of nodes of the corresponding temporal graph, and $d_{\rm m}$ quantifies thus the difference between these similarity patterns. In the Methods section, we show that we actually do not need to compute the large ($n\times n$) $M_{X_{1}}$ and $M_{X_{2}}$ matrices, but that the distance can be computed using small $d\times d$ matrices as follows:
@@
We can easily verify that $d_{\rm m}$ is a distance, as it satisfies all the requirements mentioned above and is generally not invariant under the permutation of node indices.
@@
In (3), the term $X_{1}^{T}X_{2}$ implies that $X_{1,2}$ have the same size and that their rows are ordered according to the same and known indexing of the temporal graphs’ nodes. In the case of unmatched graphs, this node correspondence is not available, and the distance, as we discussed, should also be invariant with respect to node permutations. We therefore define the distance between unmatched graphs using auxiliary vectors. Specifically, we denote with $\bm{\lambda}(A)$ the vector of ordered eigenvalues of a matrix $A$ and define our distance between two unmatched temporal graphs as
@@
where $X_{1,2}$ are, as before, the embeddings matrices of the two graphs, and $n_{1,2}=|\mathcal{V}_{1,2}|$ are their sizes. The sizes are used to normalize the covariance matrices $X_{1}^{T}X_{1}$ and $X_{2}^{T}X_{2}$ so that their eigenvalues are independent from graph size.
The covariance matrices $X_{1}^{T}X_{1}$ and $X_{2}^{T}X_{2}$ (of size $d\times d$) are indeed an appropriate argument to define the distance $d_{\rm u}$, since they are independent of the ordering of the nodes and any other orthogonal transformation applied to the columns of $X$. However, directly comparing the entries of $X_{1}^{T}X_{1}$ and $X_{2}^{T}X_{2}$ requires a one-to-one correspondence between the embedding dimensions (the columns of $X$). This mapping is unavailable because the EDRep algorithm makes no guarantees on the correspondence between the dimensions of the embedding spaces of different graphs. As a solution to this issue, we thus compare their spectra, which are invariant with respect to orthogonal transformations of the rows of $X$. This choice takes inspiration from the spectral static graph distance definitions – such as Ref. [32] – in which graphs are compared by computing the distance between the ordered sets of the eigenvalues of their matrix representations (e.g., adjacency, Laplacian, non-backtracking, etc.)

### v2

@@ Abstract
Temporal graphs are commonly used to represent time-resolved relations between entities in many natural and artificial systems. Many techniques were devised to investigate the evolution of temporal graphs by comparing their state at different time points. However, quantifying the similarity between temporal graphs as a whole is an open problem. Here, we use embeddings based on time-respecting random walks to introduce a new notion of distance between temporal graphs. This distance is well-defined for pairs of temporal graphs with different numbers of nodes and different time spans. We study the case of a matched pair of graphs, when a known relation exists between their nodes, and the case of unmatched graphs, when such a relation is unavailable and the graphs may be of different sizes. We use empirical and synthetic temporal network data to show that the distance we introduce discriminates graphs with different topological and temporal properties. We provide an efficient implementation of the distance computation suitable for large-scale temporal graphs.
@@ Introduction
Graphs are ubiquitous mathematical entities formed by a set of nodes and one of edges connecting them [1, 2, 3]. They can model a wide range of interacting systems, such as social [4, 5], technological [6], spatial [7] and biological networks [8], and owe their popularity to their ability to encode the complex relational structure of these systems. Real-world systems, moreover, often have temporal properties that cannot be encoded in static graphs and call for modeling based on time-resolved network representations, known as temporal graphs [9]. Examples of these include transportation [10] and ecological networks [11], human close-range interactions [12, 13, 14, 15, 16], collaboration networks [17, 18, 19], etc.
@@
Given their ubiquitous use in representing such diverse kinds of systems, it is crucial to be able to compare them. In fact, defining and computing a similarity measure between pairs of graphs [20] underpins many important applications and tasks, including machine learning tasks such as graph classification. However, given the high dimensionality of graphs and their structural complexity, many different notions of distance can be devised, capturing different properties of interest. Hence, many definitions of distance for static graphs were introduced, as reviewed in Refs. [21, 22, 23, 24, 25]. One of the most straightforward approaches is the edit distance [26], which counts the number of elementary changes (edge or node removal or addition) needed to transform one graph into another and only accounts for local information. Approaches that probe the global graph structure are usually based on matrix inversion [27, 28] but generally have a high computational cost and approximations are needed to achieve satisfactory computational performances. The above methods are all designed for comparing pairs of graphs with a known correspondence between nodes, a case we will refer to as matched graphs. The more general case of unmatched graphs entails defining distances for pairs of graphs for which a mapping between nodes is unavailable, and the graphs may have different numbers of nodes. In this case, one can either extract and compare a suitable set of chosen graph features, as done in Refs. [29, 30], or leverage the spectral properties of appropriate matrix representations [31, 32, 33, 34]. In this case, the computational complexity is also a bottleneck, and approximation heuristics are often required [21].
@@
In the case of temporal graphs, the temporal dimension adds to the complexity of graph comparison. Many of the distance measures mentioned above for static graphs were used to study temporal graphs [35], typically to compare the state of the graph at different points in time and to identify anomalies and change points in the temporal evolution of the graph [36, 37, 38, 39, 40]. Here instead, we study the problem of comparing temporal graphs, each taken as a whole. Unlike the snapshot-based methods mentioned above [36, 37, 38, 39, 40], we rely on a temporal embedding with a dimensionality that is independent of the temporal span of the graphs, so that we can compare temporal graphs with arbitrary, and different, temporal spans. Few notions of distance were proposed to compare temporal graphs with arbitrary time spans. In particular, the method of Ref. [41] is based on a set of features that can be tailored to a specific research question (similarly to [29] for static graphs). Ref. [42] on the other hand addresses the same problem of temporal graph comparison we tackle here, but it assumes the graphs to be matched and it is computationally inefficient (the graph dissimilarity is obtained solving an NP-hard optimization problem and the authors propose an approximation running in $\mathcal{O}(n^{3})$ operations, where $n$ is the number of graph nodes). Finally, Ref. [43] only considers the unmatched case and focuses on a specific property of temporal graphs, namely the comparison of temporal paths statistics, to provide a definition of dissimilarity.
@@
Here, we propose a computationally efficient method to compute an actual distance metric between pairs of temporal graphs, considering both the matched and unmatched cases. We build our comparison on top of graph embeddings, leveraging ideas from the local and global approaches mentioned above and encoding differences at the topological and temporal levels. Commented Python code that implements the approach described here is available at github.com/lorenzodallamico/G-DynaDist.
@@
Before delving into the details of our main contribution, we first state some basic definitions.
@@
A temporal graph $\mathcal{G}(\mathcal{V},\mathcal{E},\mathcal{W})$ is a tuple $(\mathcal{V},\mathcal{E},\mathcal{W})$, where $\mathcal{V}$ is a set of $n$ nodes, $\mathcal{E}$ a set of temporal edges between nodes, and $\mathcal{W}$ a set of edge weights. For a discrete variable $t$ denoting time, all temporal edges $e\in\mathcal{E}$ have the form $e=(i,j,t)$, representing an interaction between nodes $i$ and $j$ at time $t$ and $w_{e}$ indicates the positive weight of edge $e$. We say an edge $(ij)$ is active at time $t$ if $(i,j,t)\in\mathcal{E}$. A node $i$ is active at time $t$ if it belongs to an active edge at $t$.
The graph $\mathcal{G}(\mathcal{V},\mathcal{E},\mathcal{W})$ can be represented with a sequence of weighted adjacency matrices, $\{W_{t}\}_{t=1,\dots,T}$, where $T$ is the number of time points. Each of these matrices has size $n\times n$ and entries $(W_{t})_{ij}=w_{(i,j,t)}>0$ if $(i,j,t)\in\mathcal{E}$ and $(W_{t})_{ij}=0$ otherwise. In the following, we will refer to these matrices as “snapshots” of the temporal graph at given times and to $T$ as the number of snapshots [44]. In a more general description of temporal networks, each interaction may be represented with a tuple $(i,j,t,\tau)$ where $\tau\in\mathbb{R}^{+}$ is an interaction duration. We adopt the former notation because it is simpler to handle in the following. The two notations are completely interchangeable whenever the time is discrete and a finite temporal resolution $t_{\rm res}$ is set. In fact, all contacts with a duration within $t_{\rm res}$ are considered to be instantaneous, and $(i,j,t,\tau)$ can be represented by a set of interactions of unitary duration: $\{(i,j,t+\delta)\}_{\delta\in\{0,\dots,\tau-1\}}$. For this reason, Definition 1 is sufficiently general, as it can also describe interactions lasting more than one temporal unit. For increasing values of $t_{\rm res}$, the temporal graph gets progressively more aggregated and loses information on the temporal ordering of contacts within $t_{\rm res}$, but partial information can be retained with edge weights expressing the fraction of $t_{\rm res}$ in which the edge was active.
@@
Let $\mathcal{G}_{1}(\mathcal{V}_{1},\mathcal{E}_{1},\mathcal{W}_{1})$ and $\mathcal{G}_{2}(\mathcal{V}_{2},\mathcal{E}_{2},\mathcal{W}_{2})$ be two temporal graphs with $n_{1}$ and $n_{2}$ nodes, respectively. We say that $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are matched if $n_{1}=n_{2}$ and there exists a known bijective function $\pi:\mathcal{V}_{1}\to\mathcal{V}_{2}$ that maps each node of $\mathcal{G}_{1}$ to a node of $\mathcal{G}_{2}$. The two graphs are otherwise said to be unmatched.
When two graphs have the same size, i.e. , the same number of nodes, there are always many possible mappings $\pi$ between their nodes, but for the matched approach to be valuable, a known mapping associated with an external notion of node identity across the two graphs is necessary (e.g., nodes corresponding to the same persons in two temporal social networks). An implicit consequence of Definition 2 is that all graph pairs with a different number of nodes are considered to be unmatched. We now state the requirements for our distance measure and illustrate our main result.
Figure 1: Distance computation between two temporal graphs. The inputs to the distance function are two temporal graphs $\mathcal{G}_{1}$ (top, orange nodes) and $\mathcal{G}_{2}$ (bottom, green nodes), generally with a different number of snapshots ($T_{1}=5$, $T_{2}=4$, in the example) and a different number of nodes $n$ (here $n_{1}=6$, $n_{2}=5$). Each graph is represented as a matrix $P\in\mathbb{R}^{n\times n}$, according to (1), with entry $(ij)$ encoding the probability that a random walker goes from $i$ to $j$ following a time-respecting path (the walker’s position at each time point is indicated in blue). The matrices $P_{1},P_{2}$ are then embedded using the EDRep algorithm, mapping them to $X_{1}\in\mathbb{R}^{n_{1}\times d}$ and $X_{2}=\mathbb{R}^{n_{2}\times d}$. Finally, the matched – (3) – and unmatched – (4) – distances are calculated. Notice that a necessary condition to compute $d_{\rm m}$ is that $n_{1}=n_{2}$.
@@
In the following, we define two types of distance functions: $d_{\rm m},d_{\rm u}$ between matched and unmatched graphs, respectively. The inputs to these functions are two temporal graphs, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, with potentially different numbers of time points (snapshots) $T_{1}$ and $T_{2}$. In both cases, the distances should fulfill the following standard requirements [45]: i) non-negativity; ii) separation axiom; iii) symmetry; iv) triangle inequality. In addition, both distances should be able by design to detect differences induced by the permutation of time indices. The matched distance should also differentiate between temporal networks differing only by a permutation of node indices, while the unmatched one should be invariant under such transformation.
Our proposed distances are computed in two steps: i) we generate a temporal graph embedding, given by a matrix of size $(n\times d)$ (with $d$ independent of $T$), that encodes relevant topological and temporal properties by leveraging time-respecting random walks on the temporal network; ii) we define the distance in the embedded space, treating separately the matched and the unmatched cases. The distance computation pipeline is illustrated in Fig. 1. The steps outlined above are described in detail in the next two Sections titled Temporal graph embedding and Defining an embedding-based distance.
@@
Graphs are rich data representations, making it challenging to define simple mathematical operations to manipulate them. Graph embeddings provide an interesting avenue for doing so, as they yield a representation in vector space that preserves some of the graphs’ relevant properties [46, 47, 48, 49]. In the case of temporal graphs, the additional challenge of dealing with their temporal dimension has been tackled in many different ways: Ref. [50] represents time as an additional dimension of an adjacency tensor; in evolutionary spectral clustering [51, 52, 53, 54, 55], a different embedding is obtained for each snapshot, and smoothness conditions are implemented on their temporal evolution; more recent work [56, 57, 58] introduced approaches based on time-respecting random walks, which encode temporal properties relevant for dynamical processes occurring on temporal graphs; Refs. [39, 40] also build one embedding per snapshot for studying the graph evolution and comparing different time points. The authors argue that a graph comparison metric should depend on the whole graph structure and thus define graph instead of node embeddings.
@@
With a similar approach to [58], we start from time-respecting random walks to generate temporal graph embeddings, which we use to define a distance based on global temporal graph properties. The embeddings are obtained as the solution to an optimization problem. We define a loss function whose argument is a transition matrix $P$ that is the limiting distribution of time-respecting random walks over the temporal graph $\mathcal{G}$ to be embedded. $T$ being the number of snapshots of $\mathcal{G}$, we consider an ensemble of random walks of length $\ell$ drawn uniformly at random between $1$ and $T$, and starting from a randomly chosen node at time $t=T-\ell+1$; at each time step, the walker, situated, e.g., on node $i$, can either move to a random neighbor of $i$ in the snapshot at time $t$ or stay in place on $i$; if the walker is situated on a node that is not active at time $t$, i.e. , which has no neighbors in this snapshot, the walker stays in place with probability one. The graph snapshots on the left of Fig. 1 show examples of such time-respecting random walks on the input graphs.
@@
Specifically, given a temporal graph $\mathcal{G}$, with $W_{t}$ the snapshot adjacency matrices at times $t=1,\cdots,T$, we define for each snapshot $\hat{L}_{t}=(D_{t}+I_{n})^{-1}(W_{t}+I_{n})$, where $D_{t}$ is the degree matrix $D_{t}={\rm diag}(W_{t}\bm{1}_{n})$ of the snapshot, $I_{n}$ is the identity matrix of size $n\times n$, and $\bm{1}_{n}$ is the $n$-sized unity vector. We then define the matrix $P$ as
$P=\frac{1}{T}\sum_{\tau=1}^{T}\hat{L}_{\tau}\hat{L}_{\tau+1}\cdots\hat{L}_{T}\,\,,$ (1)
The matrices $\hat{L}_{t}$ and $P$ are sometimes referred to as temporal transition matrix and global transition matrix, respectively. The properties of time-respecting random walks depend on several properties of the temporal graph, that may include, for instance, the presence of a non-Markovian behavior, and of a broad distribution of the edge activity and inter-event durations [59]. The limiting probability matrix $P$ of time respecting random walks encodes the aforementioned properties and it is sensitive to the permutation of time indices. We remark that the random walks we use can be of arbitrary length. Hence, the matrix $P$ is, in principle, sensitive to all time scales of the temporal graph at hand.
@@
Given the above probability matrix $P$ and a chosen embedding dimensionality $d$, we now generate, for each node $i$ of $\mathcal{G}$, a unitary node embedding vector $\bm{x}_{i}\in\mathbb{R}^{d}$. To this end, we define $X\in\mathbb{R}^{n\times d}$ as a matrix where each row represents an embedding vector. We then express the loss function as:
@@
This loss function is the sum of the cross entropies between the distributions in the rows of $P$ and the variational distributions in the rows of $Q(X)$, with the addition of a regularization term. Since $Z_{i}$ is computed in $\mathcal{O}(n)$ operations for each $i$, the function $\mathcal{L}_{P}$ requires $\mathcal{O}(n^{2})$ operations to be computed. We adopt the recently proposed EDRep algorithm [60] that introduces an efficient approximation of $Z_{i}$, optimizing $\mathcal{L}_{P}$ in $O(n)$ operations. For further details, we refer the reader to the Methods section.
The matrix $X$ obtained by optimizing $\mathcal{L}_{P}$ defines the embedding of the temporal graph $\mathcal{G}$. In the following, we show how to leverage this embedding to define a distance between temporal graphs.
Defining an embedding-based distance
We want to define a distance between temporal graph embeddings satisfying the requirements stated above. Notice that the embedding vectors $\bm{x}$ are defined up to an orthogonal transformation applied to the rows of $X$. In fact, suppose $R\in\mathbb{R}^{d\times d}$ is an orthogonal matrix, and $\tilde{\bm{x}}_{i}=R\bm{x}_{i}$, then $\bm{x}_{i}^{T}\bm{x}_{j}=\bm{\tilde{x}}_{i}^{T}\bm{\tilde{x}}_{j}$ and the loss function of (2) is $\mathcal{L}_{P}(\tilde{X})=\mathcal{L}_{P}(X)$. Hence, our distance must be invariant under orthogonal transformations. Moreover, as discussed above, the distance for unmatched graphs must also be invariant with respect to node permutations.
Given these requirements, we now discuss separately our choices for the the matched and unmatched cases. In the Empirical evaluation section, we show the ability of our distances to discriminate temporal graph classes that differ for topological or temporal properties.
@@
Let $X_{1}$ and $X_{2}\in\mathbb{R}^{n\times d}$ be the embedding matrices of two matched temporal graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. To satisfy the invariance requirements discussed above, rather than defining our distance directly in terms of the matrices $X_{1}$ and $X_{2}$, we use the auxiliary matrices $M_{X_{1}}=X_{1}X_{1}^{T}$ and $M_{X_{2}}=X_{2}X_{2}^{T}$ of size $n\times n$. These matrices are by construction invariant with respect to orthogonal transformations of the embedding space and encode the global structure of the corresponding temporal graphs, as the embeddings are generated using the complete temporal network information. We define the distance $d_{\rm m}$ between matched temporal graphs as:
@@
where $\|\cdot\|_{\rm F}$ is the Frobenius norm. Each matrix $M$ encodes the similarity in the embedding space between all pairs of nodes of the corresponding temporal graph, and $d_{\rm m}$ quantifies thus the difference between these similarity patterns. In the Methods section, we show that we do not need to actually compute the large ($n\times n$) $M_{X_{1}}$ and $M_{X_{2}}$ matrices, but that the distance can be computed using the generally smaller $d\times d$ matrices as follows:
@@
We can easily verify that $d_{\rm m}$ is a distance, as it satisfies all the requirements mentioned above and is generally not invariant under the permutation of node indices. Notice that, even if $d_{\rm m}$ can be computed for any pair of graphs with the same size, this distance is only meaningful for matched graphs (i.e., with known matching). In (3), the term $X_{1}^{T}X_{2}$ implies that $X_{1}$ and $X_{2}$ have the same size and that their rows are ordered according to the same node index.
@@
In the case of unmatched graphs, there is no correspondence between nodes, and the distance, as we discussed, should also be invariant with respect to node permutations. We therefore define the distance between unmatched graphs using auxiliary vectors. Specifically, we denote with $\bm{\lambda}(A)$ the vector of ordered eigenvalues of a matrix $A$ and define our distance between two unmatched temporal graphs as
@@
where $X_{1}$ and $X_{2}$ are, as before, the embeddings matrices of the two graphs, with sizes $n_{1}=|\mathcal{V}_{1}|$ and $n_{2}=|\mathcal{V}_{2}|$. The normalization of the covariance matrices $X_{1}^{T}X_{1}$ and $X_{2}^{T}X_{2}$ is chosen so that their eigenvalues are independent from the graph size.
In the Methods section we show that the covariance matrices $X_{1}^{T}X_{1}$ and $X_{2}^{T}X_{2}$ (of size $d\times d$) are indeed an appropriate argument of the distance function $d_{\rm u}$, since they are independent of the ordering of the nodes and any other orthogonal transformation applied to the columns of $X_{1}$ and $X_{2}$. However, a direct comparison of the entries of $X_{1}^{T}X_{1}$ and $X_{2}^{T}X_{2}$ would require a one-to-one correspondence between the embedding dimensions, i.e., the columns of $X_{1}$ and $X_{2}$. This mapping is unavailable because the EDRep algorithm, like most vector-space embedding techniques, makes no guarantees on the correspondence between the dimensions of the embedding vectors corresponding to different graphs. To tackle this issue, we thus compare the spectra, which are invariant with respect to orthogonal transformations of the rows of the embedding matrices. This choice takes inspiration from spectral distance definitions for static graphs, such as Ref. [33], in which graphs are compared by computing the distance between the ordered sets of the eigenvalues of their matrix representations (e.g., adjacency, Laplacian, non-backtracking, etc.)